# Magnetic phases of the periodic Anderson model in two dimensions

### v1

@@ III.1 Half-filled case
Several comments are in order: to discuss the predictions of mean-field theory for this filling and to place our findings in the context of previous studies. We find that, in the mean-field description, the vertical stripe state – comprising ferromagnetic lines arranged in an antiferromagnetic pattern – has the lowest energy when the $f$ level is occupied. However, as outlined above, no signatures of such states were observed in the iPEPS calculations. Consequently, we did not compare the properties of this state with the iPEPS results in Fig. 5. An interesting feature of the mean-field calculation is the presence of a diagonal AF stripe-like solution, although its energy is much higher than that of the vertical stripe state. This discrepancy arises because mean-field theory can lower the energy of the Hubbard term only by inducing magnetic order; it cannot simultaneously suppress double occupancy and maintain paramagnetic behavior at specific sites. As shown in Fig. 3(a), the strong fluctuations along the $[11]$ diagonal suppress magnetic moments and enhance local correlations. Regarding energetics, we note that although mean-field theory is not a variational approach, the raw iPEPS energies (without extrapolation) are significantly lower than the mean-field ones for both $n=2$ and $n=1.5$.
@@
We also comment on previous variational Monte Carlo studies [14] and constrained path quantum Monte Carlo simulations [20], which predicted a ferromagnetic ground state and a resonating spin-density-wave type ground state, respectively. The discrepancy with our findings in the former case likely stems from the fact that only conventional antiferromagnetic, paramagnetic, and uniform ferromagnetic states were considered in the variational Monte Carlo approach. The newly identified candidate states from our work cannot be captured by these Ansätze, highlighting the need for more general trial wave functions in future studies. The results of the constrained path quantum Monte Carlo simulations, on the other hand, show some similarities with our findings. They reported a ground state characterized by the absence of onsite magnetization, strong antiferromagnetic spin correlations between diagonal sites, and weak correlations between nearest-neighbor sites on a $6\times 6$ lattice. This state could be a precursor of the long-range ordered diagonal AF state presented in this paper, although no evidence of its ferromagnetic counterpart was observed in that study.
@@ IV Conclusions
We investigated the ground-state properties of the periodic Anderson model (PAM) on the square lattice at various fillings. Using the iPEPS technique, we obtained unbiased results across a wide range of parameters. First, we examined the symmetric half-filled model and observed that increasing the hybridization leads to the disappearance of antiferromagnetic order. Beyond a critical hybridization value, local Kondo singlets are formed. Our estimate for the critical point is in excellent agreement with quantum Monte Carlo results, underscoring the accuracy of the iPEPS approach. Next, we studied the system away from the symmetric point by varying the $f$-level energy. We found that antiferromagnetic order persists in the mixed-valence regime but vanishes when the $f$-level occupancy falls below $n_{f}\sim 0.5$. A comparison with mean-field theory revealed that, while the mean-field approach provides a qualitatively correct picture, it significantly overestimates the critical point and conduction electron magnetization. This discrepancy becomes even more pronounced in the mixed-valence region.
In the second part of the paper, we focused on the doped case with $n=1.5$ electrons per site. We demonstrated the emergence of novel diagonal (anti-) ferromagnetic stripe states, characterized by strong local correlations along one diagonal. In contrast, the mean-field calculation incorrectly predicts a vertical stripe state. We also highlighted the competition between these states and concluded, based on our energy extrapolations, that the antiferromagnetic stripe state is energetically favored. Overall, our results indicate that the ground state at $n=1.5$ filling is more complex than those previously proposed in variational Monte Carlo calculations, warranting future studies with more sophisticated trial wave functions.
As an outlook, we note that we did not explore the possibility of a superconducting ground state in this model, which could provide insights into heavy-fermion superconductors. Another promising extension involves including Hund’s coupling via the Kanamori Hamiltonian to investigate how quantum fluctuations modify the phase diagram [47]. Such studies are challenging for quantum Monte Carlo methods, but as we have shown, iPEPS is capable of providing reliable results for models with large local Hilbert spaces. These problems are therefore within the reach of the iPEPS technique.
@@ Acknowledgements.
The author thanks Ö. Legeza, L. Oroszlány and P. Vancsó for useful comments and careful reading of the manuscript. This project was supported by the Hungarian National Research, Development and Innovation Office (NKFIH) through Grants No. FK142985 and No. K134983, by the Quantum Information National Laboratory of Hungary, as well as by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. We acknowledge the use of the Noctua2 cluster at the Paderborn Center for Parallel Computing (PC2). Some of the data presented here was produced using the SyTen toolkit [48, 49].

### v2

@@ III.3 Discussion
Several comments are in order: to discuss the predictions of mean-field theory for this filling and to place our findings in the context of previous studies. We find that, in the mean-field description, the vertical stripe state – comprising ferromagnetic lines arranged in an antiferromagnetic pattern – has the lowest energy when the $f$ level is occupied. However, as outlined above, no signatures of such states were observed in the iPEPS calculations. Consequently, we did not compare the properties of this state with the iPEPS results in Fig. 5. An interesting feature of the mean-field calculation is the presence of a diagonal AF stripe-like solution, although its energy is much higher than that of the vertical stripe state. This discrepancy arises because mean-field theory can lower the energy of the Hubbard term only by inducing magnetic order; it cannot simultaneously suppress double occupancy and maintain paramagnetic behavior at specific sites. As shown in Fig. 3(a), the strong fluctuations along the $[11]$ diagonal suppress magnetic moments and enhance local correlations. Regarding energetics, we note that although mean-field theory is not a variational approach, even the raw iPEPS energies (without extrapolation) are significantly lower than the mean-field ones for both $n=2$ and $n=1.5$.
@@
It is also worth putting these findings in context with previous results in the literature. Mean-field theory on the cubic lattice [16], DMFT on hypercubic lattice [49], variational Monte Carlo as well as [14] constrained path quantum Monte Carlo simulations [22] on the square lattice addressed the ground-state properties of the PAM at $n=1.5$ filling. Mean-field theory, DMFT and variational Monte Carlo predict a ferromagnetic ground state for a wide range of paramaters and occupied $f$-level. The discrepancy with our findings likely stems from the fact that only conventional antiferromagnetic, paramagnetic, and uniform ferromagnetic states were considered in the variational Monte Carlo and mean-field approaches. The newly identified candidate states from our work cannot be captured by these Ansätze, highlighting the need for more general trial wave functions or unit cells in future studies. The results of the constrained path quantum Monte Carlo simulations [22], on the other hand, show some similarities with our findings. They found that a resonating spin-density-wave ground state emerges due to the nesting property of the Fermi surface with the wave vectors $\bm{q}=(\pi,0),(0,\pi)$. This state is characterized by the absence of onsite magnetization, strong antiferromagnetic spin correlations between diagonal sites, and weak correlations between nearest-neighbor sites on a $6\times 6$ lattice. Since larger system sizes were out of reach for this technique, the question whether this state evolves to long-range order or not in the thermodynamic limit remains inconclusive. The resonating spin-density-wave ground state observed in this finite lattice might be a precursor of the long-range ordered diagonal AF state presented in this paper, although no evidence of its ferromagnetic counterpart was observed in that study.
@@ IV Conclusions
We investigated the ground-state properties of the periodic Anderson model (PAM) on the square lattice at various fillings. Using the iPEPS technique, we obtained unbiased results across a wide range of parameters. First, to benchmark the applicability of the iPEPS approach, we examined the symmetric half-filled model and observed that increasing the hybridization leads to the disappearance of antiferromagnetic order. Beyond a critical hybridization value, local Kondo singlets are formed. Our estimate for the critical point is in excellent agreement with quantum Monte Carlo results, underscoring the accuracy of the iPEPS approach. Next, we studied the system away from the symmetric point by varying the $f$-level energy. We found that antiferromagnetic order persists in the mixed-valence regime but vanishes when the $f$-level occupancy falls below $n_{f}\sim 0.5$. A comparison with mean-field theory revealed that, while the mean-field approach provides a qualitatively correct picture, it significantly overestimates the critical point and conduction electron magnetization. This discrepancy becomes even more pronounced in the mixed-valence region.
In the second part of the paper, we focused on the doped case with $n=1.5$ electrons per site. We demonstrated the emergence of novel diagonal (anti-) ferromagnetic stripe states, characterized by strong local correlations along one diagonal. We also highlighted the competition between these states and concluded, based on our energy extrapolations, that the antiferromagnetic stripe state is energetically favored. In contrast, the mean-field calculation incorrectly predicts a vertical stripe state. Overall, our results indicate that the ground state at $n=1.5$ filling is more complex than those previously proposed in variational Monte Carlo calculations, warranting future studies with more sophisticated trial wave functions.
As an outlook, we mention that it would be interesting to examine how stable the diagonal stripe phase is against doping in the vicinity of $n=1.5$ filling. Furthermore, we did not explore the possibility of a superconducting ground state in this model, which could provide insights into heavy-fermion superconductors. Another promising extension involves including Hund’s coupling via the Kanamori Hamiltonian to investigate how quantum fluctuations modify the phase diagram [51]. Such studies are challenging for quantum Monte Carlo methods, but as we have shown, iPEPS is capable of providing reliable results for models with large local Hilbert spaces. These problems are therefore within the reach of the iPEPS technique.
@@ Acknowledgements.
The author thanks Ö. Legeza, L. Oroszlány and P. Vancsó for useful comments and careful reading of the manuscript. This project was supported by the Hungarian National Research, Development and Innovation Office (NKFIH) through Grants No. FK142985 and No. K134983, by the Quantum Information National Laboratory of Hungary, as well as by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. We acknowledge the use of the Noctua2 cluster at the Paderborn Center for Parallel Computing (PC2). Some of the data presented here was produced using the SyTen toolkit [52, 53]. The numerical data that support the findings of this study are available from the author upon reasonable request.
Data availability
The data that support the findings of this article are openly available [54].
Appendix: Overview of the iPEPS algorithm
Here we give a brief overview and further technical details about the iPEPS method employed in the paper. For an in-depth review of the method we refer to Refs. [29, 30, 31, 32, 33, 34, 42], here we summarize its main aspects. As we outlined in the main text, the iPEPS on a square lattice consists of five-legged tensors. Four of them are auxiliary indices of dimension $D$ that link it to its nearest neighbors, along with a physical index of dimension $d$ representing the local Hilbert space, which is $d=16$ in our model. The bond dimension $D$ governs the accuracy of the variational Ansatz. A ground-state iPEPS simulation unfolds in two main steps. First, the tensor network is optimized to faithfully approximate the ground state of a chosen Hamiltonian. We achieve this optimization via the imaginary-time evolution. Once optimized, the iPEPS enables the calculation of physical properties – for instance, expectation values of observables. Both stages generally require contracting the two-dimensional tensor network. We discuss these stages in more detail in the following.
Contraction of the tensor network
First, we overview the contraction of an iPEPS state, as this is also a prerequisite for the different optimization techniques. The exact contraction of an iPEPS state is exponentially hard, therefore approximate but controlled contraction schemes were developed. We use the corner transfer matrix (CTM) technique [44, 45], whose main idea is shown schematically in Fig. 6 for the simplest case, where the iPEPS has a single-site unit cell.
Figure 6: Illustration of the contraction of an iPEPS with bond dimension $D$ using the CTM technique for the calculation of the norm. First, we contract the double-layer iPEPS along the physical indices, then the network of reduced tensors (with bond dimension $D^{2}$) can be approximated by a set of corner and edge tensors (with bond dimension $\chi$) that encode the information of the infinite environment.
When calculating the norm or the expectation value of a local operator, we need to contract the double-layer network shown in Fig. 6. The CTM technique makes this contraction feasible by introducing eight fixed-point tensors surrounding the iPEPS tensor in the middle. The accuracy of this technique can be controlled by the bond dimension of the environment tensors, $\chi$. The numerical cost of the CTM technique scales as $\mathcal{O}(\chi^{3}D^{6})$ [42], which is $\mathcal{O}(D^{12})$ for our choice with $\chi=D^{2}$.
Optimization of an iPEPS
We use the imaginary time evolution for the ground-state search of a given Hamiltonian, $H$. The optimization starts usually from a random initial state $|\phi\rangle$ and for long enough times
$|\Psi_{\rm GS}\rangle=\lim_{\beta\rightarrow\infty}\frac{e^{-\beta H}|\phi\rangle}{||e^{-\beta H}|\phi\rangle||}$ (A7)
the ground state $|\Psi_{\rm GS}\rangle$ is projected out. In order to make this tractable, the time evolution operator is decomposed into two-site gates using the Trotter-Suzuki formula (second order):
$e^{-\beta H}=e^{-\tau\sum_{b}{H_{b}}}=\prod_{b}e^{-\tau H_{b}}+\mathcal{O}(\tau^{2}),$ (A8)
where $\tau=\beta/N$ is the time step and the summation extends over the bonds in the supercell. The Trotter error, occurring due to the splitting of the time evolution operator to two-site gates, is well-controlled and can be reduced by making the time steps smaller. After the gate application to an iPEPS bond with bond dimension $D$, it will increase to $d^{2}D$. To keep the iPEPS state numerically feasible we need to truncate this back to $D$. Unlike matrix-product states, the iPEPS does not have a canonical form because of loops in the lattice, therefore a naive usage of singular value decomposition for the truncation produces suboptimal results. Consequently, a different strategy is utilized to find the truncated tensors. Let us consider a bond $b$ hosting the $A$ and $B$ tensors. After the gate application the new state $|\Psi^{\prime}_{A^{\prime}B^{\prime}}\rangle=e^{-\tau H_{b}}|\Psi_{AB}\rangle$ needs to be truncated to bond dimension $D$. In the full update this is achieved by finding iteratively the new $\tilde{A}$ and $\tilde{B}$ tensors with bond dimension $D$ that minimize the
$C=||\ \Psi^{\prime}_{A^{\prime}B^{\prime}}\rangle-|\Psi_{\tilde{A}\tilde{B}}\rangle||^{2}$ (A9)
cost function [33, 43]. As we mentioned in the main text the quantity $w=C/\tau$ – which is similar to the truncation error in the density-matrix renormalization group method – can be used as an error measure to extrapolate the ground-state energy to the error-free limit [46]. Since the full update requires the calculation of the whole environment in each step, this becomes computationally expensive with increasing the bond dimension. Hence, another approach, the fast-full update [43] was developed which recycles part of the CTM environment but maintains the accuracy of the update procedure at the same time. We compared both approaches for our Hamiltonian and found that they give identical results up to $D=7$, where the full update was still feasible.